OpenStudy (anonymous):

Show that (n-7)/(n+7) converges with the limit 1. When n goes from 1 to infinity.

OpenStudy (anonymous):

Does it make sense to split it up into two: \[\frac{ n }{ n+7 }\] and \[\frac{ -7 }{ n+7 }\]

ganeshie8 (ganeshie8):

good idea, just try splitting it in another way

ganeshie8 (ganeshie8):

How about : \(\dfrac{n-7}{n+7} = \dfrac{(n+7)-14}{n+7} = 1 -\dfrac{14}{n+7}\)

OpenStudy (anonymous):

Yes, that is a great idea. Now i should be able to insert it in the definition for convergence.

ganeshie8 (ganeshie8):

Yes, remember \(\lim\limits_{n\to\infty} \dfrac{1}{n} = 0\)

ganeshie8 (ganeshie8):

wait a second, are you supposed to prove this using epsilon delta definition of limit ?

OpenStudy (anonymous):

No, we have only learned the definition for convergence, and cauchy.

ganeshie8 (ganeshie8):

then we're good i hope :)

OpenStudy (anonymous):

I insert it with the limit:\[|1-\frac{ 14 }{ n+7 }-1|=|-\frac{ 14 }{ n+7 }|=\frac{ 14 }{ n+7 }\] Now this needs to be less or equal to epislon. Can I do the following: \[\frac{ 14 }{ n+7 }\le \frac{ 1 }{ n }\le \epsilon\]

OpenStudy (anonymous):

Because if i get it to \[\frac{ 1 }{ n }\], then i can isolate n: \[n \ge \frac{ 1 }{ \epsilon }\] And get that if a natural number N is bigger than \[\frac{ 1 }{ \epsilon }\] then the function is convergent for all \[n \ge N\]

ganeshie8 (ganeshie8):

so you do want to prove it using epsilong delta definition !

ganeshie8 (ganeshie8):

\[\frac{ 14 }{ n+7 }\le \frac{ 1 }{ n }\le \epsilon\] this is wrong, try again.

OpenStudy (anonymous):

Ohh, that is epsilon delta definition. I didnt know that, all my courses and stuff is on another language.

OpenStudy (anonymous):

Yea, i realize that, is there anyway that I can get to \[\frac{ 1 }{ n }\]

ganeshie8 (ganeshie8):

dont get attached to \(\dfrac{1}{n}\) so much

ganeshie8 (ganeshie8):

you cuold try somehting like this : \[\frac{ 14 }{ n+7 }\lt \frac{ 14 }{ n }\lt \epsilon\] so we want \(n\gt \dfrac{14}{\epsilon}\)

ganeshie8 (ganeshie8):

Now you're ready to start the proof

ganeshie8 (ganeshie8):

what we did so far is just the scratch work, its not proof ok

ganeshie8 (ganeshie8):

your proof must start with the value of \(N\) that works

OpenStudy (anonymous):

Ohh, i thought that was the proof.

OpenStudy (anonymous):

Yea, but \[n>14/epsilon\] shows that for any number \[N>14/epsilon\], then \[|1-\frac{ 14 }{ n+7 }-1|\le \epsilon\] be true for all \[n >= N\]

OpenStudy (anonymous):

I'm pretty sure, thats how they do in my books.

ganeshie8 (ganeshie8):

you're right, but you need to put that in proper order in your proof

ganeshie8 (ganeshie8):

if psble, may i see your complete proof ?

OpenStudy (anonymous):

Well, the problem is that it is in Danish, and I am not great at english and dont know how to make a proof in english

ganeshie8 (ganeshie8):

thats okay :) please take a look at 2nd example here : http://www.millersville.edu/~bikenaga/math-proof/limits-at-infinity/limits-at-infinity.html

ganeshie8 (ganeshie8):

your final proof must have that structure

OpenStudy (anonymous):

I will look at it, thanks

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

At the start my book lets epsilon > 0, and then they do the "scratch work", and then they kinda concludes that it is convergent.

OpenStudy (anonymous):

The only thing I havent seen / I dont understand is where they set M=max()

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

GM

OpenStudy (anonymous):

@ganeshie8 When they use M, is that the N\[N \le n\] that I have been using?

ganeshie8 (ganeshie8):

Yes M is same as N

OpenStudy (anonymous):

How old are yall

ganeshie8 (ganeshie8):

you may simply let \(M = \dfrac{14}{\epsilon}\) in our case forget about \(M = \max(0, \dfrac{14}{\epsilon})\) if it confuses you now..

OpenStudy (anonymous):

How old are yall

ganeshie8 (ganeshie8):

for our probloem, it shouldn't matter either ways

OpenStudy (anonymous):

Okay, thank you soo much! You are great :)

OpenStudy (anonymous):

Okay I let y'all go

OpenStudy (anonymous):

Bye